Samples of 4 cards are drawn from a population of 6 cards numbered 1-6.



Construct a sampling distribution of the sample means and answer the following questions



1. How many sample of size 4 can be drawn from the population?


2.What are the possible means?


3.What is the probability of getting 4 as a mean?


4.What is the probability of getting 3.5 as a mean?

1
Expert's answer
2022-05-16T16:41:42-0400

We have population values 1,2,3,4,5,6, population size N=6 and sample size n=4.

Mean of population "(\mu)" = "\dfrac{1+2+3+4+5+6}{6}=3.5"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{1}{6}(6.25+2.25+0.25"


"+0.25+2.25+6.25)=\dfrac{17.5}{6}"

"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{17.5}{6}}\approx1.707825"

1. Select a random sample of size 4 without replacement. We have a sample distribution of sample mean.

The number of possible samples which can be drawn without replacement is "^{N}C_n=^{6}C_4=15."


2.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 1,2,3,4 & 10/4 \\
 \hdashline
 2 & 1,2,3,5 & 11/4 \\
 \hdashline
 3 & 1,2,3,6 & 12/4 \\
 \hdashline
 4 & 1,2,4,5 & 12/4 \\
 \hdashline
 5 & 1,2,4,6 & 13/4 \\
 \hdashline
 6 & 1,2,5,6 & 14/4 \\
 \hdashline
 7 & 1,3,4,5 & 13/4 \\
 \hdashline
 8 & 1,3,4,6 & 14/4 \\
 \hdashline
 9 & 1,3,5,6 & 15/4 \\
 \hdashline
 10 & 1,4,5,6 & 16/4 \\
 \hdashline
 11 & 2,3,4,5 & 14/4 \\
 \hdashline 
 12 & 2,3,4,6 & 15/4 \\
 \hdashline 
 13 & 2,3,5,6 & 16/4 \\
 \hdashline 
 14 & 2,4,5,6 & 17/4\\
 \hdashline 
 15 & 3,4,5,6 & 18/4 \\
 \hdashline 
 
\end{array}"




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 10/4 & 1/15 & 10/60 & 100/240 \\
 \hdashline
 11/4 & 1/15 & 11/60 & 121/240 \\
 \hdashline
 12/4 & 2/15 & 24/60 & 288/240 \\
 \hdashline
 13/4 & 2/15 & 26/60 & 338/240 \\
 \hdashline
 14/4 & 3/15 & 42/60 & 588/240 \\
 \hdashline
 15/4 & 2/15 & 30/60 & 450/240 \\
 \hdashline
 16/4 & 2/15 & 32/60 & 512/240 \\
 \hdashline
 17/4 & 1/15 & 17/60 & 289/240 \\
 \hdashline
 18/4 & 1/15 & 18/60 & 324/240 \\
 \hdashline
\end{array}"



3.


"P(\bar{X}=4)=\dfrac{2}{15}"

4.


"P(\bar{X}=3.5)=\dfrac{3}{15}=\dfrac{1}{5}"